HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 57     Entry time: Tue Apr 1 23:56:18 2025
Author: Shahab, Carlo, Zac, Yue, Michael 
Category: DAQ 
Subject: Change of RSA01 settings for new harmonic=212. 
23:55 stop the RSAs. 

410:
last file from previous setting (harmonic=211): 2025.04.01.23.55.21.712.
new plunger position for 410 cavity: 33.5 inside and outside.
new center frequency: 410535000 Hz.

245:
harmonic didn't change, still 126.
adjusted center frequency to 243997000 Hz.

00:31 restart the RSAs.
new files from: 2025.04.02.00.31.11.796.
ELOG V3.1.5-fc6679b